The Diver Time company makes digital watches for scuba divers. The company models its profits with the function P(d) = 2d^2 + 13d -6 , where d is the number of watches sold, in hundreds, and P(d) is the company’s profit, in tens of thousands of dollars.
a. how much profit is made when a 100,000 watches are sold?
b. how many watches must the company sell in order to break even?

Found 2 solutions by MowMow, ikleyn:
Answer by MowMow(42)

P(d) = 2d^2 + 13d -6
a. how much profit is made when a 100,000 watches are sold?
P(100,000) = 2x10^4x10^4 + 13x10^4 -6 = $200,129,994 profit.

Answer by ikleyn(52779)

.
The Diver Time company makes digital watches for scuba divers. The company models its profits with the function
P(d) = 2d^2 + 13d -6 , where d is the number of watches sold, in hundreds, and P(d) is the company’s profit,
in tens of thousands of dollars.
a. how much profit is made when a 100,000 watches are sold?
b. how many watches must the company sell in order to break even?
~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution and calculations by @MowMow in his post for part  (a)  are  INCORRECT.

        I came to bring a correct solution.


We are given that 100,000 watches are sold.  100,000 is  1000  hundreds,
so, we should substitute 1000 for 'd' into the formula 

    P(1000) = 2*1000^2 + 13*1000 - 6 = 2012994.


So, the profit is  2012994  in tens of thousands of dollars, or  20,129,940,000  dollars.


ANSWER to part (a).  The profit is 20,129,940,000 dollars.

Solved correctly.